# Volichenko-type metasymmetry of braided Majorana qubits

### v1

@@ 11 Conclusions
The mixed-bracket Heisenberg-Lie algebras, which have been introduced in Section 8 and reproduce the multi-particle parastatistics of the braided Majorana qubits, differ from the quonic “mixed brackets” as derived in formulas (239,240). Indeed, the (93) building blocks $A_{k;N}^{\dagger}$ of the creation operators of the $N$-particle sector and their $A_{k;N}=(A_{k;N}^{\dagger})^{\dagger}$ annihilation counterparts do not satisfy the quonic algebra (239) for any nontrivial $q\neq\pm 1$ value; it is easily shown that they satisfy, for any given $k=1,2,\ldots,N$, an ordinary anticommutator:
@@
Formulas (266,267,268) reproduce, for $g=\frac{1}{3}$, the corresponding “minimal” formulas presented in (90,91,92).
@@
The matrices in (264,265) reproduce the mixed brackets of, respectively, the $2$-particle and $3$-particle Heisenberg-Lie algebras at the level $s=3$ (that is, the third root of unity). They induce the same algebras as the ones recovered from the corresponding minimal representations introduced in Section 5 (see, e.g., formula (5) specialized at the $g=\frac{1}{3}$ value).

### v2

@@ 9 Color parafermionic oscillators from the $s\rightarrow\infty$ limit
9.1 The $s\rightarrow\infty$ limit versus the bosonic multi-particle sectors
It is worth discussing the difference between the $s\rightarrow\infty$ limit versus the bosonic construction of the multi-particle sectors of the quantized model.
At first one should recall that the $2\times 2$ matrices $\alpha,\beta,\gamma,\delta$ introduced in (14) close the ${\mathfrak{gl}}(1|1)$ superalgebra with $\beta,\gamma$ odd (fermionic) elements.
In an alternative description the three matrices $\beta,\gamma,\alpha-\delta$ can also be regarded as the generators of the ${\mathfrak{sl}}(2)$ algebra in the fundamental representation. This description induces a bosonic quantization of the model.
The $N=2,3,4,\ldots$ multi-particle sectors can be constructed for both bosonic and level-$s$ fermionic versions. In all cases the corresponding multi-particle Hilbert spaces ${\cal H}_{N}$ are subspaces of ${\cal H}^{\otimes N}$, the tensor product of the two-dimensional, single-particle Hilbert space ${\cal H}$; one gets, see (23), ${\cal H}_{N}\subset{\cal H}^{\otimes N}$. The dimension of ${\cal H}^{\otimes N}$ is $2^{N}$.
By taking into account that the finite-dimensional representation theory of the noncompact group $SL(2,{\mathbb{R}})$ is equivalent to the representation theory of $SU(2)$, the bosonic formulation gives a nice interpretation of ${\cal H}^{\otimes N}$ as a decomposition into irreducible representations of the spin-$\frac{1}{2}$ reps tensors:
${\cal H}^{\otimes N=2}\equiv[1/2]\otimes[1/2]=[0]\oplus[1]\Rightarrow{\textrm{dim }}4=1+3,$
${\cal H}^{\otimes N=3}\equiv[1/2]\otimes[1/2]\otimes[1/2]=[1/2]\oplus([1/2]\oplus[3/2])\Rightarrow{\textrm{dim }}8=2\times 2+1\times 4$
[…]
The dimensions of the fermionic (at level-$s$) and of the bosonic multi-particle Hilbert spaces can be read from the coproducts applied to the creation operator $\gamma$.
For ordinary $(s=2)$ fermions the (3) $t=1$ choice implements, as discussed in Section 3, the Pauli exclusion principle: only two states (a vacuum state and an excited fermionic state) survive, for any given $N$, in the multi-particle sector. Therefore, for ordinary $s=2$ fermions the dimensions of the multi-particle Hilbert spaces are
${\textrm{dim}}({\cal H}_{N=1})={\textrm{dim}}({\cal H}_{N=2})={\textrm{dim}}({\cal H}_{N=3})=\ldots=2.$ (234)
Unlike the level-$s$ fermionic versions which use the $t$-dependent braided tensor products (16), the bosonic multi-particle quantization induced by ${\cal U}({\mathfrak{sl}}(2))$ is obtained from an ordinary (not braided) tensor product. In particular, the following relation is satisfied:
$({\mathbb{I}}_{2}\otimes\gamma)\cdot(\gamma\otimes{\mathbb{I}}_{2})=(\gamma\otimes{\mathbb{I}}_{2})\cdot({\mathbb{I}}_{2}\otimes\gamma).$ (235)
A straightforward combinatorics shows that the dimensions of the multi-particle bosonic Hilbert spaces ${\cal H}_{bos;N}\subset{\cal H}^{\otimes N}$ are given by
${\textrm{dim}}({\cal H}_{bos;N})=N+1.$ (236)
The bosonic coproduct selects, within ${\cal H}^{\otimes N}$, the highest spin subspace in the (9.1) irreducible decompositions. One gets
${\textrm{dim}}({\cal H}_{bos;N=2})\equiv[1],\quad{\textrm{dim}}({\cal H}_{bos;N=3})\equiv[\frac{3}{2}]\quad{\textrm{and, in general,}}\quad{\textrm{dim}}({\cal H}_{bos;N})\equiv[\frac{N}{2}].$
It is important to notice that, for any $N$, the dimension of the bosonic $N$-particle Hilbert space coincides with the dimension of the fermionic untruncated $N$-particle Hilbert space recovered in the $s\rightarrow\infty$ limit. This feature, on the other hand, does not imply that these two $N$-particle Hilbert spaces should be identified.
The reason is that, even in the $s\rightarrow\infty$ limit, the untruncated multi-particle fermionic Hilbert space keeps track of the odd-graded nature of $\gamma$. This is why, in that limit, parafermionic oscillators and not bosonic oscillators are recovered. It was already mentioned, at the beginning of this Section, that the ${\mathbb{Z}}_{2}$-grading of the Hilbert space implies a superselection of the states (bosons versus fermions) and that states of different gradings cannot be superposed. This superselection rule is absent in the purely bosonic theory.
This dual description (bosonic versus ${\mathbb{Z}}_{2}$-graded) is a common feature in quantum models, even encountered in the quantization of the harmonic oscillator. The results of the famous [31] 1950 Wigner’s paper ”Do the Equations of Motion Determine the Quantum Mechanical Commutation Relations?” can be reinterpreted in modern language, see [32], in this framework. Basically, in [31] Wigner flips the statistics of the $a^{\dagger},a$ creation/annihilation operators of the harmonic oscillator, inducing an ${\mathfrak{osp}}(1|2)$ spectrum-generating superalgebra; he recovers the spectrum of the harmonic oscillator from a lowest weight rep of ${\mathfrak{osp}}(1|2)$, instead of the standard Fock-vacuum derivation. In 1950 Lie superalgebras had not yet been introduced in mathematics and, possibly, Wigner’s construction can be the first appearance of what is now called the ${\mathfrak{osp}}(1|2)$ superalgebra.
In the ${\mathfrak{osp}}(1|2)$ derivation of the spectrum of the harmonic oscillator, the “bosonic” versus “fermionic” labellings of the states refer to their respective parities under the $x\leftrightarrow-x$ parity transformation of the space coordinate; the vacuum state (proportional to the gaussian $e^{-\frac{1}{2}x^{2}}$ for a convenient choice of parameters) is bosonic (even), the first excited state (proportional to $xe^{-\frac{1}{2}x^{2}}$) is fermionic (odd) and so on.
@@ 11 Conclusions
The mixed-bracket Heisenberg-Lie algebras, which have been introduced in Section 8 and reproduce the multi-particle parastatistics of the braided Majorana qubits, differ from the quonic “mixed brackets” as derived in formulas (244,245). Indeed, the (93) building blocks $A_{k;N}^{\dagger}$ of the creation operators of the $N$-particle sector and their $A_{k;N}=(A_{k;N}^{\dagger})^{\dagger}$ annihilation counterparts do not satisfy the quonic algebra (244) for any nontrivial $q\neq\pm 1$ value; it is easily shown that they satisfy, for any given $k=1,2,\ldots,N$, an ordinary anticommutator:
@@
Formulas (271,272,273) reproduce, for $g=\frac{1}{3}$, the corresponding “minimal” formulas presented in (90,91,92).
@@
The matrices in (269,270) reproduce the mixed brackets of, respectively, the $2$-particle and $3$-particle Heisenberg-Lie algebras at the level $s=3$ (that is, the third root of unity). They induce the same algebras as the ones recovered from the corresponding minimal representations introduced in Section 5 (see, e.g., formula (5) specialized at the $g=\frac{1}{3}$ value).